you and two friends each randomlypick a movie from 7 choices. what is the probibility that you will all pick the same movie?
Answer by stanbon(75887)

you and two friends each randomlypick a movie from 7 choices. what is the probibility that you will all pick the same movie?
----
Binomial Problem with n = 3 and p(same) = 1/7
----
P(x = 3) = 3C3*(1/7)^3*(6/7)^0 = (1/7)^3 = 0.0029
===============
Cheers,
Stan H.
===============